EXTENSIONS AND CORRECTIONS FOR: “A CONVEX GEOMETRIC APPROACH TO COUNTING THE ROOTS OF A

POLYNOMIAL SYSTEM”

JOSEPH MAURICE ROJAS

Abstract. This brief note corrects some errors in the paper quoted in the title, highlights a combinatorial result which may have been overlooked, and points to further improvements in recent literature.